Risk free rate = 0.03 Market rate = 0.11 N is the current year. In general, the EPS and dividend for N will be given.

3M has a beta of 0.95, an expected dividend of $5.82 next year, and an 8% growth rate. If it’s trading at $201.10, is it over or under valued?
Answer #1

First we calculate the Required rate of return(R), then we get the current price of the stock based on D1=$5.82 with growth(g) 8%.

So, R = rf + B (rm - rf) = 0.03 + 0.95 (0.11 - 0.03) = 0.106

Current expected Price of Stock = D1 / (R - g) = 5.82 / (0.106 - 0.08) = $223.85

The Stock is under-valued as the prevailing price is $201.10.